# Incommensurate Twisted Bilayer Graphene: emerging quasi-periodicity and stability

### v1

@@ D.2 Constraints on the resonant terms
    Using that $W_{aa}(k_{0},k)=^{d}-W_{aa}(-k_{0},k)$ we get $W_{aa}(0,p_{F}^{\omega})=\partial_{1}W_{aa}(0,p_{F}^{\omega})=\partial_{2}W_{aa}(0,p_{F}^{\omega})=0$. Similarly, $W_{bb}(0,p_{F}^{\omega})=\partial_{1}W_{bb}(0,p_{F}^{\omega})=\partial_{2}W_{bb}(0,p_{F}^{\omega})=0$.
@@
    From $W_{ab}(k_{0},k)=^{b}W_{ba}(k_{0},-k)=^{a}W^{*}_{ba}(-k_{0},k)$ we get $W_{ab}(0,p_{F}^{\omega})=W^{*}_{ba}(0,p_{F}^{\omega})$, $\partial_{1}W_{ab}(0,p_{F}^{\omega})=\partial_{1}W^{*}_{ba}(0,p_{F}^{\omega})$, $\partial_{2}W_{ab}(0,p_{F}^{\omega})=\partial_{2}W^{*}_{ba}(0,p_{F}^{\omega})$. Moreover $W_{ab}(k_{0},k)=^{d}W_{ab}(-k_{0},k)$ hence $\partial_{0}W_{ab}(0,p_{F}^{\omega})=0$.
@@
    $\partial_{0}W_{aa}(k_{0},k)=^{a}-\partial_{0}W_{aa}^{*}(-k_{0},-k)=^{b}-\partial_{0}W_{aa}^{*}(-k_{0},k)$ hence is pure imaginary at $k_{0}=0$; moreover $\partial_{0}W_{aa}(k_{0},k)=^{c}\partial_{0}W_{bb}(k_{0},-k)=^{a}-\partial_{0}W^{*}_{bb}(-k_{0},k)$ so that $\partial_{0}W_{aa}(0,p_{F}^{\omega})=\partial W_{bb}(0,p_{F}^{\omega})=iz$ with $z$ real

### v2

@@ D.2 Constraints on the resonant terms
    Using that $W_{aa}(k_{0},k)=^{d}-W_{aa}(-k_{0},k)$ we get $W_{aa}(0,K_{\omega})=\partial_{1}W_{aa}(0,K_{\omega})=\partial_{2}W_{aa}(0,K_{\omega})=0$. Similarly, $W_{bb}(0,K_{\omega})=\partial_{1}W_{bb}(0,K_{\omega})=\partial_{2}W_{bb}(0,K_{\omega})=0$.
@@
    From $W_{ab}(k_{0},k)=^{b}W_{ba}(k_{0},-k)=^{a}W^{*}_{ba}(-k_{0},k)$ we get $W_{ab}(0,K_{\omega})=W^{*}_{ba}(0,K_{\omega})$, $\partial_{1}W_{ab}(0,K_{\omega})=\partial_{1}W^{*}_{ba}(0,K_{\omega})$, $\partial_{2}W_{ab}(0,K_{\omega})=\partial_{2}W^{*}_{ba}(0,K_{\omega})$. Moreover $W_{ab}(k_{0},k)=^{d}W_{ab}(-k_{0},k)$ hence $\partial_{0}W_{ab}(0,K_{\omega})=0$.
@@
    $\partial_{0}W_{aa}(k_{0},k)=^{a}-\partial_{0}W_{aa}^{*}(-k_{0},-k)=^{b}-\partial_{0}W_{aa}^{*}(-k_{0},k)$ hence is pure imaginary at $k_{0}=0$; moreover $\partial_{0}W_{aa}(k_{0},k)=^{c}\partial_{0}W_{bb}(k_{0},-k)=^{a}-\partial_{0}W^{*}_{bb}(-k_{0},k)$ so that $\partial_{0}W_{aa}(0,K_{\omega})=\partial W_{bb}(0,K_{\omega})=iz$ with $z$ real